OpenStudy (anonymous):

Katie Ward is building a children's playhouse. She wants each window to have an area of 315 square inches for adequate air. For eye appeal, she wants each window to be six inches higher than wide. What are the dimensions of the window?

OpenStudy (cwrw238):

x(x+ 6) = 315 solve for x ( x = width)

OpenStudy (anonymous):

How

OpenStudy (cwrw238):

x^2 + 6x = 315 x^2 + 6x - 315 = 0 (x + 21 )(x - 15 ) x must be 15 so width of window = 15 and height is 15+ 6 = 21

OpenStudy (cwrw238):

have done this stuff in algebra tana?

OpenStudy (cwrw238):

* have you done

OpenStudy (anonymous):

no not yet

OpenStudy (cwrw238):

ah right i should have asked u first the other way to do this is by factoring 315: find the prime factors of 315 3 ) 315 3 ) 105 3 ) 35 5 ) 7 so 315 = 3 x 3 x 5 x 7 now , the difference between height and width is 6 inches so from the above factors we pi ck 3 x 5 = 15 and 3 x 7 = 21 dimensions of window are w=15, h = 21